OpenStudy (anonymous):

find three consecutive even numbers so that the sum of the first and third numbers is 22 less than three times the second number

OpenStudy (lgbasallote):

let x = first even number x+2 = consecutive even number x+4 = next consecutive even number x + x+ 4 = 3(x+2) - 22 2x + 4 = 3x + 6 - 22 22 + 4 - 6 = 3x - 2x 20 = x x+2 = 22 x+ 4 = 24